OpenStudy (anonymous):

is cos(x+y)+cos(x-y)=2cosx(cosy) an identity?

terenzreignz (terenzreignz):

Find out... if you can turn the left-side into the right side using a series of conversions, then it's an identity... In English, it means that regardless of what permissible values of x and y we put, the equality holds... (compare for instance, with, 3x + 9 = 12, which is not true for all values of x, rather, only for x = 1 )

terenzreignz (terenzreignz):

Okay... we have these known identities that will help... \[\Large \cos(\color{red}x+\color{blue}y) = \cos(\color{red}x)\cos(\color{blue}y) - \sin(\color{red}x)\sin(\color{blue}y)\]\[\Large \cos(\color{red}x-\color{blue}y) = \cos(\color{red}x)\cos(\color{blue}y) + \sin(\color{red}x)\sin(\color{blue}y)\]

terenzreignz (terenzreignz):

So for the left-side, we have \[\Large \cos(\color{red}x+\color{blue}y) +\Large \cos(\color{red}x-\color{blue}y)\] Using the identities shown, these two can be expanded.